### v1

# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

Abstract
A common problem in Message-Passing Neural Networks is oversquashing – the limited ability to facilitate effective information flow between distant nodes. Oversquashing is attributed to the exponential decay in information transmission as node distances increase. This paper introduces a novel perspective to address oversquashing, leveraging properties of global and local non-dissipativity, that enable the maintenance of a constant information flow rate. Namely, we present SWAN, a uniquely parameterized model GNN with antisymmetry both in space and weight domains, as a means to obtain non-dissipativity. Our theoretical analysis asserts that by achieving these properties, SWAN offers an enhanced ability to transmit information over extended distances. Empirical evaluations on synthetic and real-world benchmarks that emphasize long-range interactions validate the theoretical understanding of SWAN, and its ability to mitigate oversquashing.
Machine Learning, ICML, oversquashing, over-squashing, deep graph network, graph neural network, long range interactions, ordinary differential equations
1 Introduction
A critical issue that limits Message-Passing Neural Networks (MPNNs) (Gilmer et al., 2017), a class of GNNs, is the oversquashing problem (Alon & Yahav, 2021; Di Giovanni et al., 2023). In the oversquashing scenario, the capacity of MPNNs to transmit information between nodes exponentially decreases as their distance increases, which imposes challenges on modeling long-range interactions, which are often necessary for real-world tasks (Dwivedi et al., 2022). At the same time, it has been shown in Haber et al. (2017); Chen et al. (2018a) that neural networks can be interpreted as the discretization of ordinary differential equations (ODEs). Building on these observations and understandings, similar concepts were utilized to forge the field of differential-equations inspired GNNs (DE-GNNs), as shown in (Poli et al., 2019; Chamberlain et al., 2021b) and subsequent works. Through this view, it is possible to design GNNs with strong inductive biases, such as smoothness (Chamberlain et al., 2021b), energy preservation (Eliasof et al., 2021; Rusch et al., 2022), node-wise non-dissipativity (Gravina et al., 2023), and more.
(a) Source
(b) Target
(c) Diffusion
(d) Local Non-
Dissipativity
(e) Local and Global Non-Dissipativity
Figure 1: An illustration of the ability of Global and Local Non-Dissipativity (e) to propagate information to distant nodes, from the source (a) to the target (b). Other dynamics, such as diffusion (c) cannot achieve the desired behavior, while Local Non-Dissipativity (d) has only limited effect. Our SWAN implements (e).
In this paper, we are interested in addressing the oversquashing problem in a principled manner, accompanied by theoretical understanding through the prism of DE-GNNs. Current literature offers several approaches to mitigate oversquashing, such as adding a virtual global node (Gilmer et al., 2017; Cai et al., 2023), graph rewiring (Gasteiger et al., 2019; Topping et al., 2022; Karhadkar et al., 2023; Nguyen et al., 2023), as well as using graph transformers (Shi et al., 2021; Dwivedi & Bresson, 2021; Rampášek et al., 2022). Some of the methods above focus on providing ad-hoc mechanisms for the network to reduce oversquashing. Instead, in this paper, we are interested in theoretically understanding and improving the information propagation capacity of the network, thereby mitigating oversquashing. Specifically, we will show that it is possible to design GNNs with a constant flow of information, unlike typical diffusion-based methods. To this end, we take inspiration from the recent ADGN (Gravina et al., 2023), a non-dissipative GNN. In ADGN, it was shown that by incorporating an antisymmetric transformation to the learned channel-mixing weights, it is possible to obtain a locally, i.e., node-wise, non-dissipative behavior. Here, we propose SWAN (Space-Weight ANtisymmetry), a novel GNN model that is both globally (i.e., graph-wise) and locally (i.e., node-wise) non-dissipative, achieved by space and weight antisymmetric parameterization. To understand the behavior of SWAN and its effectiveness in mitigating oversquashing, we propose a global, i.e., graph-wise, analysis, and show that compared to ADGN (Gravina et al., 2023), our SWAN is both globally and locally non-dissipative. The immediate implication of this property is that SWAN is guaranteed to have a constant information flow rate, thereby mitigating oversquashing. Such a property is visualized in Figure 1, where SWAN shows improved capacities of propagating information across the graph, with respect to diffusion and local non-dissipative approaches. To complement our theoretical analysis, we experiment with several long-range benchmarks, from synthetic to real-world datasets and tasks.
Main Contributions. In this paper, we present the following contributions. (i) A novel graph perspective theoretical analysis of the stability and non-dissipativity of antisymmetric DE-GNNs, providing a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNN model. (ii) We propose SWAN, a space and weight antisymmetric DE-GNN with a constant information flow rate, and an increased distant node interaction sensitivity. (iii) We experimentally verify our theoretical understanding of SWAN, on both synthetic and real-world datasets. Our experiments show score improvements of up to 117% in synthetic datasets, as well as competitive results on real-world benchmarks compared to existing methods, from MPNNs to other DE-GNNs, and graph transformers – highlighting the importance of global and local non-dissipative behavior offered by SWAN.
2 Preliminaries
We provide an overview of oversquashing problem in GNNs, followed by a brief discussion of DE-GNNs and antisymmetric weights in DE-GNNs. We then provide a mathematical background, which is thoroughly used in this paper to analyze and understand the properties of SWAN.
2.1 Oversquashing
Oversquashing (Alon & Yahav, 2021) refers to the shortcoming of a GNN when transferring information between distant nodes, and is a common problem in Message-Passing Neural Networks (MPNNs). This shortcoming typically increases in tandem with the distance between the nodes, which hampers the ability of MPNNs to model complex behaviors that require long-range interactions. Standard MPNNs update node states by considering neighborhood information. To allow a node to receive information from $k$-hops distant nodes, an MPNN must employ at least $k$ layers, otherwise, it will suffer from under-reaching, because the two nodes are too far to interact with each other. However, the stacking of multiple message-passing layers also causes each node to receive an exponentially growing amount of information, as multiple hops are considered. This exponential growth of information, combined with the finite and fixed number of channels (features), can lead to a potential loss of information. As a consequence, the long-range effectiveness of MPNN is reduced.
2.2 GNNs Inspired by Differential-Equations
Notation. We consider a graph ${\mathcal{G}=(\mathcal{V},\mathcal{E})}$ as a system of interacting entities, referred as nodes, where $\mathcal{V}$ is a set of $n$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a set of $m$ edges. We define the neighborhood of the $u$-th node as the set of nodes directly attached to it, i.e., $\mathcal{N}_{u}=\{v|(v,u)\in\mathcal{E}\}$. The $u$-th node is associated with a (possibly time-dependent) hidden feature vector $\mathbf{x}_{u}(t)\in\mathbb{R}^{d}$ with $d$ features, which provides a representation of the node at time $t$ in the system. The term $\mathbf{X}(t)=[\mathbf{x}_{0}(t),\ldots,\mathbf{x}_{n-1}(t)]^{\top}$ is an $n\times d$ matrix that represents the nodes state (features) at time $t$.
DE-GNNs. The core idea of DE-GNN architectures is to view GNNs as the discretization of the following ODE:
${\frac{\partial\mathbf{x}_{u}(t)}{\partial t}}=s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right),$ (1)
where $\mathbf{x}_{u}(t=0)={\bf x}^{(0)}$ and $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ is a spatial aggregation function that depends on the graph $\mathcal{G}$ and the node features $\mathbf{x}_{u}(t)$. Specifically, it is common to implement $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ with graph diffusion, combined with a channel mixing operator implemented by a multilayer perceptron (MLP). Some examples of such methods were proposed in (Chamberlain et al., 2021b; Eliasof et al., 2021; Thorpe et al., 2022; Choi et al., 2023), and others.
Antisymmetric Weights in DE-GNNs. Learnable antisymmetric weights were studied in ADGN (Gravina et al., 2023), whose model is summarized below for completeness:
${\frac{\partial\mathbf{x}_{u}(t)}{\partial t}}=\sigma\Bigl{(}(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)+\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})\Bigr{)},$ (2)
where ${\bf W},\ {\bf V}\in\mathbb{R}^{d\times d}$ are learnable weights, $\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})$ is any permutation invariant neighborhood aggregation function, and $\sigma$ is a non-linear activation function. The main theoretical property of ADGN (Gravina et al., 2023) is that it allows a stable and non-dissipative propagation from a local, node-wise perspective.
2.3 Mathematical Background
In this work, we are interested in studying the stability and non-dissipativity propagation of information in DE-GNNs. Therefore, we follow the analysis techniques presented in Chang et al. (2019); Gravina et al. (2023) and focus on analyzing the sensitivity of an ODE solution with respect to its initial condition, i.e.,
$\frac{d}{dt}\left(\frac{\partial\mathbf{x}(t)}{\partial\mathbf{x}(0)}\right)=\mathbf{J}(t)\frac{\partial\mathbf{x}(t)}{\partial\mathbf{x}(0)}.$ (3)
Here, we present an overview of the various outcomes that can arise from this analysis.
We follow the results and assumptions from (Chang et al., 2019; Gravina et al., 2023), and consider the Jacobian, $\mathbf{J}(t)$ in Equation (3) to not change significantly over time. In Appendix 7, we provide a discussion of the justification as well as numerical verification of our the assumption. Given that, we can apply results from autonomous differential equations (Ascher & Petzold, 1998) and solve Equation 3 analytically as follows:
$\frac{\partial\mathbf{x}(t)}{\partial\mathbf{x}(0)}=e^{t\mathbf{J}}=\mathbf{T}e^{t\mathbf{\Lambda}}\mathbf{T}^{-1}=\mathbf{T}\big{(}\sum_{k=0}^{\infty}\frac{(t\mathbf{\Lambda})^{k}}{k!}\big{)}\mathbf{T}^{-1},$ (4)
where $\mathbf{\Lambda}$ is the diagonal matrix whose non-zero entries contain the eigenvalues of $\mathbf{J}$ (i.e., $\lambda_{i}$), and $\mathbf{T}$ has the eigenvectors of $\mathbf{J}$ as columns. We observe that the qualitative behavior of $\partial\mathbf{x}(t)/\partial\mathbf{x}(0)$ is determined by the real parts of the eigenvalues of $\mathbf{J}$, leading to three different behaviors: (i) instability, (ii) dissipativity (i.e., information loss), (iii) non-dissipativity (i.e., information preservation).
Instability is observed when $\max_{i=1,...,d}Re(\lambda_{i}(\mathbf{J}))>0$. In this scenario, a small perturbation of the initial condition (i.e., a perturbation on the input graph) would cause an exponential divergence in node representations.
Dissipativity occurs when $\max_{i=1,...,d}Re(\lambda_{i}(\mathbf{J}))<0$. Only local neighborhood information is preserved by the system since the term $\partial\mathbf{x}(t)/\partial\mathbf{x}(0)$ would vanish exponentially fast over time, thereby making the nodes’ representation insensitive to differences in the input graph.
Non-dissipativity is obtained when $Re(\lambda_{i}(\mathbf{J}))=0$. In this scenario, the system is stable, and the magnitude of $\partial\mathbf{x}(t)/\partial\mathbf{x}(0)$ is constant over time, and the input graph information is effectively propagated through the successive transformations into the final nodes’ representations.
3 SWAN: Space-Weight Antisymmetric GNN
We now turn to present SWAN, space and weight antisymmetric GNN. We analyze its theoretical behavior and show that it is both global (i.e., graph-wise) and local (i.e., node-wise) non-dissipative. As a consequence, one of the key features of SWAN is that it has a constant global information flow rate, unlike common diffusion GNNs. Therefore, SWAN should theoretically be able to propagate information between any nodes with a viable path in the graph, allowing to mitigate oversquashing. Figure 2 exemplifies the differences between dissipative, local non-dissipative, and global and local non-dissipative systems.
Figure 2: The difference among non-dissipative and dissipative behaviors. With global (i.e., graph-wise) and local (i.e., node-wise) non-dissipative behavior (a), information is propagated between any pair of nodes with a viable path in the graph. Therefore, such a behavior increases the long-range effectiveness of the model, with propagation reaching distant nodes. A model exhibiting local non-dissipative behavior (b) enhances only the long-term memory capacity of individual nodes. A model demonstrating dissipative behavior (c) exhibits a convergence of node features toward non-informative values.
Space and Weight Antisymmetry. We define SWAN by including a new term that introduces antisymmetry in the aggregation function, i.e.,
$(\displaystyle\frac{\partial\mathbf{x}_{u}(t)}{\partial t}=\sigma\Bigl{(}$$(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)$
$+\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})$
$+\beta\Psi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{Z})\Bigr{)}.$ (5)
where $\mathbf{W}$, $\mathbf{V}$, and $\mathbf{Z}$ are learnable weight matrices. $\Phi$ and $\Psi$ are permutation invariant neighborhood aggregation functions, where $\Psi$ performs antisymmetric aggregation. While $\Psi$ can assume various forms of antisymmetric aggregation functions with imaginary eigenvalues, and $\Phi$ can be any aggregation function, in this paper we explore the following family of parameterizations:
$\begin{split}\Phi&=\sum_{v\in\mathcal{N}_{u}}(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top})\mathbf{x}_{v}(t),\\ \Psi&=\sum_{v\in\mathcal{N}_{u}}((\tilde{{\bf A}}_{uv}-\tilde{{\bf A}}_{vu})(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{x}_{v}(t),\end{split}$ (6)
where $\tilde{\mathbf{A}},\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ are neighborhood aggregation matrices that can be either pre-defined or learned. In our experiments, we consider two instances of Equation 6. The first, where $\tilde{\mathbf{A}},\hat{\mathbf{A}}$ are pre-defined by the random walk and symmetrically normalized adjacency matrices, respectively. The second, allows $\tilde{\mathbf{A}},\hat{\mathbf{A}}$ to be learned, as we describe later in Appendix A. More importantly, as we show now, the blueprint of $\Psi$ and $\Phi$, which is described in Equation 6 stems from a theoretical analysis of SWAN, that shows its ability to be non-dissipative both locally and globally , therefore leading to a globally constant information flow, regardless of time, that is, the model’s depth. Lastly, we note that the general formulation of $\Phi$ and $\Psi$ provide a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNNs.
3.1 Node-wise Analysis of SWAN
We reformulate Section 3 to consider the formulation of $\Phi$ and $\Psi$ as in Equation 6, reading:
$\frac{\partial\mathbf{x}_{u}(t)}{\partial t}=\sigma$ $\Bigl{(}(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)$
$+\sum_{v\in\mathcal{N}_{u}}(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top})\mathbf{x}_{v}(t)$
$+\beta\sum_{v\in\mathcal{N}_{u}}((\tilde{{\bf A}}_{uv}-\tilde{{\bf A}}_{vu})(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{x}_{v}(t)\Bigr{)}.$ (7)
SWAN is locally non-dissipative. Following the sensitivity analysis introduced in Section 2.3, we show that SWAN is stable and non-dissipative from the node perspective, i.e., it is locally non-dissipative. Therefore, we study the Jacobian matrix of our node-wise reformulation of SWAN.
In this case, the Jacobian $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$ of Section 3.1 is composed of :
$\mathbf{M}_{1}$ $=\mathrm{diag}\Bigl{[}\sigma^{\prime}\Bigl{(}(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)$
$\hskip 28.45274pt+\sum_{v\in\mathcal{N}_{u}}(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top})\mathbf{x}_{v}(t)$
$\hskip 28.45274pt+\beta\sum_{v\in\mathcal{N}_{u}}(\tilde{{\bf A}}_{uv}-\tilde{{\bf A}}_{vu})(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{x}_{v}(t)\Bigl{)}\Bigr{]},$ (8)
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})+(\hat{{\bf A}}_{uu}+\hat{{\bf A}}_{uu})(\mathbf{V}-{\bf V}^{\top}).$ (9)
Following results from (Chang et al., 2019; Gravina et al., 2023), only the eigenvalues of $\mathbf{M}_{2}$ determine the local stability and non-dissipativity of the system in Section 3 for the final behavior of the model. Specifically, if the real part of all the eigenvalues of ${\bf M}_{2}$ is zero, then stability and non-dissipativity are achieved. We note that this is indeed the case in our system, since the real part of the eigenvalues of anti-symmetric matrices is zero, and ${\bf M}_{2}$ is composed of a summation of two anti-symmetric matrices.
3.2 Graph-wise Analysis of SWAN
While the node perspective analysis is important because it shows the long-term memory capacity of individual nodes, as illustrated in Figure 2(b), it overlooks the pairwise interactions between nodes, which is equivalently described by considering the properties of Section 3 with respect to the graph, and is illustrated in Figure 2(a).
As we show below, our SWAN is a globally non-dissipative architecture, and therefore it allows the constant rate of information flow and interactions between all nodes, independently of time $t$, equivalent to the network’s depth. Therefore, we deem that SWAN’s non-dissipativity behavior is beneficial to address oversquashing in MPNNs.
SWAN is globally non-dissipative. We start by reformulating Section 3.1 from a node-wise formulation to a graph-perspective formulation, as follows:
$\frac{\partial\mathbf{X}(t)}{\partial t}=\sigma\Bigl{(}$ $\mathbf{X}(t){(\mathbf{W}-\mathbf{W}^{\top})}$
$+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-{\bf V}^{\top})}$
$+\beta{(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})}\Bigr{)}.$ (10)
We therefore turn to analyze the ODE in Section 3.2 from a graph perspective, using the sensitivity analysis introduced in Section 2.3. Applying the vectorization operator, the Jacobian $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$ writes as the multiplication of:
$\mathbf{M}_{1}$ $=\mathrm{diag}\Bigl{[}\mathrm{vec}\Bigl{(}\sigma^{\prime}\Bigl{(}\mathbf{I}\mathbf{X}(\mathbf{W}-\mathbf{W}^{\top})+$
$\hskip 28.45274pt+(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-{\bf V}^{\top})$
$\hskip 28.45274pt+\beta(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})\Bigr{)}\Bigr{)}\Bigr{]}$ (11)
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})^{\top}\otimes{\bf I}+$
$\hskip 28.45274pt+(\mathbf{V}-{\bf V}^{\top})^{\top}\otimes(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})$
$\hskip 28.45274pt+\beta(\mathbf{Z}+\mathbf{Z}^{\top})^{\top}\otimes(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top}),$ (12)
where ${\bf I}\in\mathbb{R}^{n\times n}$ is the identity matrix, $\rm{vec}$ is the vectorization operator, and $\otimes$ is the Kronecker product (see Appendix H for more details).
Similar to our node-wise analysis in Section 3.1, $\mathbf{M}_{1}$ in Section 3.2 is a diagonal matrix, and therefore to obtain stability and non-dissipativity, we need to demand that ${\bf M}_{2}$ from Section 3.2 has eigenvalues with real part equal to zero. Indeed, we see that ${\bf M}_{2}$ satisfies this condition because it is composed of a summation of three anti-symmetric matrices, whose eigenvalues are all purely imaginary, i.e., have a real part of zero. Thus, we conclude that SWAN (Section 3.2) is both stable and globally and locally non-dissipative.
We now show that the properties of global stability and non-dissipativity allow us, theoretically, to design GNNs that can mitigate oversquashing.
Theorem 3.1 (SWAN has a constant global information propagation rate).
The information propagation rate among the graph nodes $\cal{V}$ is constant, $c$, independently of time $t$:
$\left\|\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}\mathbf{X}(0))}\right\|=c,$ (13)
Proof.
Let us consider the following equation:
$\frac{d}{dt}\left(\frac{\partial\mathbf{X}(t)}{\partial\mathbf{X}(0)}\right)=\frac{d}{dt}\left(\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}\right)=\mathbf{J}(t)\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}.$ (14)
We follow the assumption in Chang et al. (2019); Gravina et al. (2023) that the Jacobian, $\mathbf{J}(t)$, does not change significantly over time, then we can apply results from autonomous differential equations and solve Equation 14:
$\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}=e^{t\mathbf{J}}=\mathbf{T}e^{t\mathbf{\Lambda}}\mathbf{T}^{-1}=\mathbf{T}\big{(}\sum_{k=0}^{\infty}\frac{(t\mathbf{\Lambda})^{k}}{k!}\big{)}\mathbf{T}^{-1},$ (15)
where $\mathbf{\Lambda}$ is the diagonal matrix whose non-zero entries contain the eigenvalues of $\mathbf{J}$, and $\mathbf{T}$ has the eigenvectors of $\mathbf{J}$ as columns, as in Section 2.3. As we have previously shown, it holds that $Re(\lambda_{i}(\mathbf{J}(t)))=0$ for $i=1,...,d$, since the Jacobian is the result of the multiplication between a diagonal matrix and an antisymmetric matrix. Therefore, the magnitude of $\partial\mathbf{X}(t)/\partial\mathbf{X}(0)$ is constant over time, and the input graph information is effectively propagated through the successive layers into the final nodes’ features. ∎
The outcome of Theorem 3.1 is that regardless of the integration time $t$ (equivalent to $\ell=t/\epsilon$ layers of SWAN, where $\epsilon$ is the step size), the information between nodes will continue to propagate at a constant rate, unlike diffusion-based GNNs that exhibit an exponential decay in the propagation rate with respect to time, as shown below.
Theorem 3.2 (Time Decaying Propagation in Diffusion GNNs).
A diffusion-based network with Jacobian eigenvalues with magnitude ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|\ ,\ i\in\{0,\ldots,n-1\}$ has an exponentially decaying information propagation rate, as follows:
$\left\|\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}\right\|=\|e^{-t\mathbf{K}}\|,$ (16)
See proof in Appendix B. This result means that with diffusion methods, after sufficiently large integration time $t$, i.e., layers, they will not be able to effectively share new information between nodes, as in early layers of the network. On the contrary, our SWAN maintains the same capacity, independent of time, meaning it retains its ability to share information across nodes with the same effectiveness in each layer of the network, regardless of its depth. We illustrate this discussion in Figure 1 and Figure 2.
3.3 The Benefit of Spatial Antisymmetry
While oversquashig was not mathematically defined in (Alon & Yahav, 2021), it was recently proposed in (Topping et al., 2022; Di Giovanni et al., 2023) to quantify the level, or lack of oversquashing, by measuring the sensitivity of node embedding after $\ell$ layers with respect to the input of another node $\left\|\partial\mathbf{x}_{v}(\ell)/\partial\mathbf{x}_{u}(0)\right\|$. Furthermore, they bounded this sensitivity score on a MPNN (Gilmer et al., 2017) as:
$\left\|\frac{\partial\mathbf{x}_{v}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|\leq\underbrace{(c_{\sigma}wp)^{\ell}}_{model}\underbrace{(\mathbf{O}^{\ell})_{vu}}_{topology}$ (17)
where $c_{\sigma}$ is the Lipschitz constant of non linearity $\sigma$, $w$ is the maximal entry-value over all weight matrices, and $p$ is the embedding dimension. The term $\mathbf{O}=c_{r}\mathbf{I}+c_{a}\mathbf{A}\in\mathbb{R}^{n\times n}$ is the message passing matrix adopted by the MPNN, with $c_{r}$ and $c_{a}$ being the weighted contributions of the residual and aggregation term, respectively. Oversquashing occurs if the right-hand side of Equation 17 is too small.
The sensitivity of SWAN. In light of the discussion above, we now provide a theoretical bound of our SWAN, with its proof in Appendix C.
Theorem 3.3 (SWAN sensitivity upper bound).
Consider SWAN (Section 3), with $\ell$ layers, and $u,v\in\mathcal{V}$ two connected nodes of the graph. The sensitivity of $v$’s embedding after $\ell$ layers with respect to the input of node $u$ is
$\left\|\frac{\partial\mathbf{x}_{v}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|\leq\underbrace{(c_{\sigma}wp)^{\ell}}_{model}\underbrace{((c_{r}\mathbf{I}+c_{a}\mathbf{A}+\beta c_{b}\mathbf{S})^{\ell})_{vu}}_{topology}$ (18)
with $c_{\sigma}$ the Lipschitz constant of non-linearity $\sigma$, $w$ is the maximal entry-value of all weight matrices, $p$ the embedding dimension, $\mathbf{A}$ the graph shift operator, $\mathbf{S}=(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})$ the antisymmetric graph operator, and $c_{r}$ and $c_{a}$ the weighted contribution of the residual term and aggregation term.
The result of Theorem 3.3 indicates that the added antisymmetric term $\Psi$ contributes to an increase in the measured upper bound. This result, together with the constant rate of information flow obtained from Theorem 3.1, holds the potential to theoretically mitigate oversquashing using SWAN.
4 Experiments
Objectives. We empirically evaluate our SWAN and compare it with literature methods. Specifically, we seek to address the following questions: (i) Can SWAN effectively propagate information to distant nodes? (ii) Can SWAN accurately predict graph properties that are related to long-range interactions? (iii) How does SWAN perform on real-world long-range benchmarks?
Baselines. In our experiments, the performance of SWAN is compared with various GNN baselines, namely: (i) MPNNs with linear complexity with respect to the size of nodes and edges, similarly to the complexity of our SWAN, i.e., GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2018), GIN (Xu et al., 2019), GINE (Hu et al., 2020), and GCNII (Ming Chen et al., 2020). (ii) DE-GNNs, as a direct comparison to the family of GNNs considered in this work, such as DGC (Wang et al., 2021), GRAND (Chamberlain et al., 2021b), GraphCON (Rusch et al., 2022), and ADGN (Gravina et al., 2023). (iii) Oversquashing designated techniques such as Graph Transformers and Higher-Order GNNs, i.e., Transformer (Vaswani et al., 2017; Dwivedi & Bresson, 2021), DIGL (Gasteiger et al., 2019), MixHop (Abu-El-Haija et al., 2019), SAN (Kreuzer et al., 2021), GraphGPS (Rampášek et al., 2022), and DRew (Gutteridge et al., 2023).
Experimental Details. We provide an extensive discussion about the datasets used in our experiments in Appendix D. We discuss the complexity of SWAN and provide runtimes in Appendix F. Our method is implemented using PyTorch (Paszke et al., 2017) and PyTorch-Geometric (Fey & Lenssen, 2019), and we will openly release the code to reproduce our empirical analysis upon acceptance.
SWAN Variants. In the following experiments, we leverage the general formulation of our method and explore two main variants of SWAN, each distinguished by the implementation of the aggregation terms $\hat{{\bf A}},\ \tilde{{\bf A}}$ in the functions $\Psi$ and $\Phi$, as shown in Equation 6. Specifically, we consider
  1. (1)
    SWAN which implements the aggregation terms using pre-defined operators, which are the symmetric normalized and random walk adjacency matrices, as described in Appendix A.2.
  2. (2)
    SWAN-learn which utilizes the learned aggregation terms described in Appendix A.2.
Both variants follow our parameterization from Equation 6, and therefore are in line with the theoretical analysis provided in Section 3, and in particular Theorem 3.1. We refer the reader to Appendix E.1 for a detailed description.
4.1 Graph Transfer
Setup. We consider three graph transfer tasks, whose goal is to transfer a label from a source node to a target node, placed at distance $k$ in the graph. Due to oversquashing and dissipative behavior of existing GNN methods that are based on diffusion, the performance is expected to degrade as $k$ increases. We initialize nodes with a random valued feature and we assign values “1” and “0” to source and target nodes, respectively. We consider three graph distributions, i.e., line, ring, crossed-ring, as illustrated in Figure 4. We consider four different distances $k=\{3,5,10,50\}$. By increasing $k$, we increase the complexity of the task and require long-range information. In Appendix D.1 and E.2 we provide additional details about the dataset and the task.
Results. Figure 3 reports the results of the transfer graph tasks. The ring-graph proved to be the hardest, leading to higher errors when compared with other graphs. Overall, baselines struggle to accurately transfer the information through the graph, especially when the distance is high, i.e., #$\#hops\geq 10$. Differently, non-dissipative methods, such as ADGN and SWAN, achieve low errors across all distances. Moreover, SWAN consistently outperforms ADGN, empirically supporting our theoretical findings that SWAN can better propagate information among distant nodes.
(a) Line                                         (b) Ring                                     (c) Crossed-Ring
Figure 3: Information transfer on (a) Line, (b) Ring, and (c) Crossed-Ring graphs. While baseline approaches struggle to transfer the information accurately, non-dissipative methods like ADGN and our SWAN achieve low errors.
4.2 Graph Property Prediction
Setup. We address tasks involving the prediction of three graph properties - Diameter, Single-Source Shortest Paths (SSSP), and node Eccentricity on synthetic graphs as introduced in (Corso et al., 2020), and follow the experimental setup outlined in (Gravina et al., 2023), and further discussed in Appendix D.2 and E.3.
Results. In Table 1 we compare SWAN and SWAN-learn with other MPNNs and DE-GNNs. Our results indicate that SWAN improves performance across all baselines. Specifically, SWAN-learn yields the lowest (best) results across all models, showing an improvement of up to 117% with respect to the runner up model. We note that similarly to the transfer task in Section 4.1, solving Graph Property Prediction tasks necessitates capturing long-term dependencies. Hence, successful prediction requires the mitigation of oversquashing. For instance, in the eccentricity task, the goal is to calculate the maximal shortest path between a node $u$ and all other nodes. Thus, it is crucial to propagate information not only from the direct neighborhood of $u$ but also from nodes that are considerably distant.
Table 1: Mean test set $log_{10}(\mathrm{MSE})$ and std averaged over 4 random weight initializations on the Graph Property Prediction tasks. The lower, the better. First, second, third and fourth best results for each task are color-coded.
Model Diameter SSSP Eccentricity
MPNNs
GCN 0.7424±0.0466 0.9499±${}_{\pm 9.18\cdot 10^{-5}}$ 0.8468±0.0028
GAT 0.8221±0.0752 0.6951±0.1499 0.7909±0.0222
GraphSAGE 0.8645±0.0401 0.2863±0.1843 0.7863±0.0207
GIN 0.6131±0.0990 -0.5408±0.4193 0.9504±0.0007
GCNII 0.5287±0.0570 -1.1329±0.0135 0.7640±0.0355
DE-GNNs
DGC 0.6028±0.0050 -0.1483±0.0231 0.8261±0.0032
GRAND 0.6715±0.0490 -0.0942±0.3897 0.6602±0.1393
GraphCON 0.0964±0.0620 -1.3836±0.0092 0.6833±0.0074
ADGN -0.5188±0.1812 -3.2417±0.0751 0.4296±0.1003
Ours
SWAN -0.5249±0.0155 -3.2370±0.0834 0.4094±0.0764
SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.219
4.3 Long-Range Graph Benchmark
Setup. We follow the settings in Dwivedi et al. (2022) and test SWAN on the Peptides-func and Peptides-struct datasets from the Long-Range Graph Benchmark (LRGB), on which we elaborate in Appendix D.3 and E.4. As metrics, we consider the average precision (AP) for the Peptides-func, and mean-absolute-error (MAE) for Peptides-struct.
[…]
Table 2: Performance of standard MPNNs, graph Transformers, DE-GNNs, and our SWAN across two LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded.
Model Peptides- Peptides- Pascal
func struct voc-sp
AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$
MPNNs
GCN 0.5930±0.0023 0.3496±0.0013 0.1268±0.0060
GINE 0.5498±0.0079 0.3547±0.0045 0.1265±0.0076
GCNII 0.5543±0.0078 0.3471±0.0010 0.1698±0.0080
GatedGCN 0.5864±0.0077 0.3420±0.0013 0.2873±0.0219
Transformers
Transformer+LapPE 0.6326±0.0126 0.2529±0.0016 0.2694±0.0098
SAN+LapPE 0.6384±0.0121 0.2683±0.0043 0.3230±0.0039
GraphGPS+LapPE 0.6535±0.0041 0.2500±0.0005 0.3748±0.0109
DE-GNNs
GRAND 0.5789±0.0062 0.3418±0.0015 0.1918±0.0097
GraphCON 0.6022±0.0068 0.2778±0.0018 0.2108±0.0091
ADGN 0.5975±0.0044 0.2874±0.0021 0.2349±0.0054
Ours
SWAN 0.6313±0.0046 0.2571±0.0018 0.2796±0.0048
SWAN-learn 0.6751±0.0039 0.2485±0.0009 0.3192±0.0250
4.4 Ablation Study
In this section, we study the different components of SWAN, namely, the contribution of guaranteed global (i.e., graph-wise) and local (i.e., node-wise) non-dissipativity, and the benefit the spatial antisymmetry.
Table 3: Performance of different versions of SWAN on Graph Property Prediction and LRGB tasks. Results are averaged over 4 random weight initializations on the Graph Property Prediction, while over 3 on the LRGB. The first, second, and third best results for each task are color-coded.
Model Diam. SSSP Ecc. Peptides- Peptides-
func struct
$log_{10}$(MSE)$\downarrow$ $log_{10}$(MSE)$\downarrow$ $log_{10}$(MSE)$\downarrow$ AP $\uparrow$ MAE $\downarrow$
Weight Antisymmetry Only
SWANβ=0 -0.3882±0.0610 -3.2061±0.0416 0.5573±0.0247 0.6195±0.0067 0.2703±0.0023
Bounded Non-Dissipative
SWAN-ne -0.5497±0.0766 -3.1913±0.0762 0.3792±0.1514 0.6119±0.0037 0.2672±0.0012
SWAN-ne-learn -0.5631±0.0694 -3.5296±0.0831 0.1317±0.1253 0.6249±0.0051 0.2606±0.0007
Global and Local Non-Dissipative
SWAN -0.5249±0.0155 -3.2370±0.0834 0.4094±0.0764 0.6313±0.0046 0.2571±0.0018
SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.219 0.6751±0.0039 0.2485±0.0009
The importance of global and local non-dissipativity. To verify the contribution of the global and local non-dissipativity offered by SWAN, we also evaluate the performance of the following variants, which can deviate from being globally and locally non-dissipative, although in a bounded manner, as we discuss in Appendix E.1. Specifically, we consider the non-enforced (NE) variants of SWAN and SWAN-learn, which differ from the original version by the use of unconstrained weight matrix $\mathbf{V}$, rather than forcing it to be antisymmetric. These two additional variants are called SWAN-ne and SWAN-learn-ne, respectively.
Table 3 presents the performance of several SWAN variants on the Graph Property Prediction and LRGB tasks. The best performance on both synthetic and real-world problems is achieved through a global and local non-dissipative behavior by SWAN-learn, showcasing the importance of global and local non-dissipativity.
The benefit of spatial antisymmetry. Table 3 highlights the advantages of the spatial antisymmetry in Section 3. By setting $\beta=0$, we obtain a model with antisymmetry solely in the weight space, exhibiting only a local non-dissipative behavior. Our results showcase a noteworthy performance improvement when spatial antisymmetry is employed. This is further supported by the improved performance of -ne versions of SWAN, which do not guarantee both global and local non-dissipative behavior, compared to SWANβ=0. Thus, spatial antisymmetry emerges as a beneficial factor in effective information propagation.
5 Related Work
Graph Neural Networks based on Differential Equations. Adopting the interpretation of convolutional neural networks (CNNs) as discretization of ODEs and PDEs (Ruthotto & Haber, 2020; Chen et al., 2018b; Zhang et al., 2019) to GNNs, works like GCDE (Poli et al., 2019), GODE (Zhuang et al., 2020), GRAND (Chamberlain et al., 2021b), PDE-GCND (Eliasof et al., 2021), GRAND++ (Thorpe et al., 2022), DGC (Wang et al., 2021), and others, propose to view GNN layers as time steps in the integration of the heat equation. This perspective allows to control the diffusion (smoothing) in the network and understand the problem of oversmoothing (Nt & Maehara, 2019; Oono & Suzuki, 2020; Cai & Wang, 2020) in GNNs. Differently, Choromanski et al. (2022) propose an architecture with an attention mechanism based on the heat diffusion kernel. Other architectures like PDE-GCNM (Eliasof et al., 2021) and GraphCON (Rusch et al., 2022) propose to mix diffusion and oscillatory processes as a feature energy preservation mechanism. Other recent works have continued to explore this family of GNNs, by proposing additional mechanisms such as anti-symmetry (Gravina et al., 2023), reaction-diffusion-based dynamics (Wang et al., 2022; Choi et al., 2023), (Zhao et al., 2023; Eliasof et al., 2023) that propose models convection-diffusion and advection-reaction-diffusion to allow directed information transportation, and (Maskey et al., 2023) that formalize a fractional Laplacian ODE. While the aforementioned works, focus on the spatial aggregation term of DE-GNNs, as well as our SWAN, the temporal domain of DE-GNNs has also been studied recently in Eliasof et al. (2024); Kang et al. (2024). For a comprehensive review of differential equations and dynamical systems based GNN, we refer the interested reader to Han et al. (2023).
Oversquashing in MPNNs. The issue of oversquashing in MPNNs has been widely recognized due to its impact on the inability of GNNs to effectively transfer information across nodes that are far apart in the graph (Alon & Yahav, 2021). In recent years, several strategies to alleviate oversquashing have been studied. One class of approaches involves graph rewiring, where methods such as SDRF (Topping et al., 2022) modify the original edge set to densify the graph as a preprocessing step. Similarly, GRAND (Chamberlain et al., 2021b), BLEND (Chamberlain et al., 2021a), and DRew (Gutteridge et al., 2023) dynamically adjust graph connectivity based on updated node features. Transformer-based approaches (Shi et al., 2021; Dwivedi & Bresson, 2021; Ying et al., 2021; Wu et al., 2023) opt to sidestep oversquashing by enabling message passing between all pairs of nodes. An alternative approach is the incorporation of non-local dynamics to construct dense graph communication functions, as seen in methods like FLODE (Maskey et al., 2023), which uses the fractional power of the graph shift operator. Similarly, QDC (Markovich, 2023) employs a dense quantum diffusion kernel, and G2TN (Toth et al., 2022) captures the diffusion path to mitigate oversquashing. Despite the success of these techniques in addressing oversquashing, a potential drawback is the increased complexity associated with propagating information at each update, often linked to denser graph shift operators. We also refer the reader to a recent survey in Shi et al. (2023).
Non-dissipative Systems. Non-dissipative Systems are characterized by the absence of energy dissipation, thus playing a crucial role in various domains such as physics (Goldstein et al., 2002), engineering (Ogata, 2010), and machine learning. In the context of the latter, non-dissipative systems are utilized to model information flow efficiently. The effectiveness of a non-dissipative behavior for enhancing the computational power of reservoir computing and recurrent neural networks, has been demonstrated in Gallicchio et al. (2017) and Chang et al. (2019), respectively. In the context of GNNs, it was shown in Gravina et al. (2023) that local non-dissipativity is beneficial for long interaction modeling.
6 Summary
In this work, we have presented SWAN (Space-Weight ANtisymmetry), a novel Differential Equation GNN (DE-GNN) designed to address the oversquashing problem. SWAN incorporates both global (i.e., graph-wise) and local (i.e., node-wise) non-dissipative properties through space and weight antisymmetric parameterization, and provides a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNN. Our theoretical and experimental results emphasize the significance of global and local non-dissipativity, achieved by SWAN. For such reasons, we believe SWAN represents a significant step forward in addressing oversquashing in GNNs.
Appendix A Architectural Details
A.1 Integration of SWAN
The ODE that defines SWAN follows the general DE-GNNs form, presented and discussed in Section 2.2. While there are various ways to integrate these equations (see for example various integration techniques in Ascher & Petzold (1998)), we follow the common forward Euler discretization approach, which is abundantly used in DE-GNNs literature (Chamberlain et al., 2021b; Eliasof et al., 2021; Gravina et al., 2023; Rusch et al., 2022). Formally, using the forward Euler method to discretize SWAN (Section 3.2) yields the following graph neural layer:
${\bf X}^{(\ell+1)}={\bf X}^{(\ell)}+\epsilon\sigma\Bigl{(}\mathbf{X}^{(\ell)}{(\mathbf{W}-\mathbf{W}^{\top})}\\ +{\Phi(\mathbf{A},\mathbf{X}^{(\ell)},\mathbf{V})}\\ +\beta{\Psi(\mathbf{A},\mathbf{X}^{(\ell)},\mathbf{Z})}\Bigr{)}.$
Note, that in this procedure we replace the notion of time with layers, and therefore instead of using $t$ to denote time, we use $\ell$ to denote the step or layer number. Here, $\epsilon$ is the discretization time steps which replaces the infinitesimal $dt$ from Section 3.2. Furthermore, we note that similar to other DE-GNNs (Chamberlain et al., 2021b; Gravina et al., 2023) we use weight-sharing, such that the learnable matrices ${\bf W},{\bf V},{\bf Z}$, as well as the spatial aggregation matrices $\hat{{\bf A}},\tilde{{\bf A}}$ in $\Phi,\Psi$ (as shown in Equation 6, and discussed in A.2) are shared across layers.
Lastly, we notice that Euler’s forward method is stable if $(1+\epsilon\lambda(\mathbf{J}(t)))$ resides within the unit circle in the complex plane for all eigenvalues of the system (Ascher & Petzold, 1998). However, given that the Jacobian matrix’s eigenvalues are purely imaginary with our method, it implies $|1+\epsilon\lambda(\mathbf{J}(t))|>1$. Consequently, Equations (3.1) and (3.2) become unstable when solved using the forward Euler’s method. To strengthen the stability of the numerical discretization method, similar to Gravina et al. (2023) we introduce a small positive constant $\gamma>0$, which is subtracted from the diagonal elements of the weight matrix $\mathbf{W}$, with the aim of placing back $(1+\epsilon\lambda(\mathbf{J}(t)))$ within the unit circle.
A.2 Spatial Aggregation Terms
In Equation 6 we utilize two aggregation terms, denoted by $\hat{{\bf A}}$ and $\tilde{{\bf A}}$. The first, $\tilde{{\bf A}}$, is used to populate the weight antisymmetric term $\Phi$, while the second, $\hat{{\bf A}}$, is used to populate the space antisymmetric $\Psi$. As discussed, in our experiments we consider two possible parameterizations of these terms, on which we now elaborate.
Pre-defined $\hat{{\bf A}},\ \tilde{{\bf A}}$. In this case we denote our architecture as SWAN and utilize the symmetric normalized adjacency matrix and the random walk normalized adjacency matrix for $\hat{{\bf A}},\ \tilde{{\bf A}}$, respectively. Formally:
$\hat{{\bf A}}={\bf D}^{-1/2}{\bf A}{\bf D}^{-1/2},\quad\tilde{{\bf A}}={\bf D}^{-1}{\bf A},$ (19)
where ${\bf A}$ is the standard binary adjacency matrix that is induced by the graph connectivity $\mathcal{E}$, and ${\bf D}$ is the degree matrix of the graph. We observe that the implementation of $\hat{{\bf A}}$ and $\tilde{{\bf A}}$ can be treated as a hyperparameter. To show this, in our experiments, we consider both the symmetric normalized adjacency matrix and the original adjacency matrix as implementations of $\hat{{\bf A}}$.
Learnable $\hat{{\bf A}},\ \tilde{{\bf A}}$. Here, we denote our architecture by SWAN-learn, and we use a multilayer perceptron (MLP) to learn edge-weights according to the original graph connectivity, to implement learnable $\hat{{\bf A}},\ \tilde{{\bf A}}$. Specifically, we first define edge features as the concatenation of the initial embedding of input node features ${\bf X}^{(0)}$ of neighboring edges. Formally, the edge features of the $(u,v)\in\cal{E}$ edge, read:
$f_{(u,v)\in\cal{E}}^{in}=x_{u}^{(0)}\oplus x_{v}^{(0)},\ f_{(u,v)\in\cal{E}}^{in}\in\mathbb{R}^{2d},$ (20)
where $\oplus$ denotes the channel-wise concatenation operator. Then, we embed those features using a 2 layer MLP:
$f_{(u,v)\in\cal{E}}^{emb}={\rm{ReLU}}({\bf K}_{2}\sigma({\bf K}_{1}(f_{((u,v)\in\cal{E})}^{in}))),$ (21)
where ${\bf K}_{1}\in\mathbb{R}^{d\times 2d}$ and ${\bf K}_{2}\in\mathbb{R}^{n\times n}$ are learnable linear layer weights, and $\sigma$ is an activation function which is a hyperparameter of our method described in E.5. By averaging the feature dimension and gathering the averaged edge features $f_{(u,v)\in\cal{E}}^{emb}$ into a sparse matrix ${\bf F}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, such that ${\bf F}_{u,v}=\frac{1}{d}\sum f_{(u,v)\in\cal{E}}^{emb}$ we define the learned spatial aggregation terms as follows:
$\hat{{\bf A}}_{{\bf F}}={\bf D}_{{\bf F}}^{-1/2}{\bf F}{\bf D}_{{\bf F}}^{-1/2},\quad\tilde{{\bf A}}_{{\bf F}}={\bf D}_{{\bf F}}^{-1}{\bf F},$ (22)
where ${\bf D}_{{\bf F}}$ is the degree matrix of ${\bf F}$, i.e., a matrix with the column sum of ${\bf F}$ on its diagonal and zeros elsewhere.
A.3 The role of $\gamma$ and $\epsilon$
The $\gamma$ term in SWAN refers to a small positive constant subtracted from the diagonal elements of the weight matrix to balance the stability of the numerical discretization method (i.e., Euler’s forward method). According to Ascher & Petzold (1998), the Euler’s forward method applied to Section 3 is stable when $|1+\epsilon\lambda(\mathbf{J}(t))|>1$, with $\epsilon$ the step size. Since SWAN has $Re(\lambda(\mathbf{J}(t)))=0$, subtracting the constant $\gamma$ ensures that the Euler method remains stable. Therefore, as $\gamma$ approaches 0, the discretization method becomes increasingly unstable, requiring taking a smaller step size $\epsilon$.
A.4 Applicability of SWAN to general MPNNs
Nowadays, most GNNs rely on the concepts introduced by the Message Passing Neural Network (MPNN) (Gilmer et al., 2017), which is a general framework based on the message passing paradigm. A general MPNN updates the representation for a node $u$ by using message and update functions. The first (message function) is responsible for defining the messages between nodes and their neighbors. On the other hand, the update function has the role of collecting (aggregating) messages and updating the node representation. Our SWAN, in Section 3, operates according to the MPNN paradigm, with the functions $\Phi$ and $\Psi$ that operate as the message function, while the sum operator among the node and neighborhood representations is the update function. Therefore, SWAN can be interpreted as a special case of an MPNN with the aformentioned parameterization, and can potentially be applied to different types of MPNNs.
Appendix B Proof of Theorem 3.2
The following proof follows the sensitivity analysis introduced in Section 2.3.
Proof.
Let us assume a diffusion-based network whose Jacobian’s eigenvalues are represented by the diagonal matrix $\mathbf{\Lambda}$, and let denote the eigenvalues magnitude by a diagonal matrix ${\bf K}\in\mathbb{R}_{+}^{n\times n}$ such that ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|$ for $i\in\{0,\ldots,n-1\}$. Applying the vectorization operator, it is true that :
$\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}=e^{t\mathbf{J}}$ (23)
As it is known from Evans (1998), diffusion-based network are characterized by Jacobian’s eigenvalues with negative real part. Indeed, diffusion DE-GNNs are based on the heat equation. Therefore, the right-hand side of the ODE is the graph Laplacian, reading $\frac{\partial{\bf X}(t)}{\partial t}=-{\bf L}{\bf X}(t)$. Therefore, following our derivations in Section 2.3, we analyze $-{\bf L}$. It is known that the graph Laplacian has non-negative eigenvalues, and therefore $-{\bf L}$ has non-positive eigenvalues. Thus, the Jacobian has non-positive eigenvalues. Assuming a connected graph (i.e., it exists at least one edge), then there is at least one value that is strictly negative in $-{\bf L}$, and any entry that is not non-negative will be equal to zero. Therefore, we can write ${\bf J}=-{\bf K}$, leading to the equation
$\left\|\frac{\partial\rm{vec}(\mathbf{X}(t))}{\partial\rm{vec}(\mathbf{X}(0))}\right\|=\|e^{-t{\bf K}}\|.$ (24)
Therefore, the information propagation rate among the graph nodes exponentially decays over time. Note, that at $t\rightarrow\infty$, it will converge to 1, which is exponentially lower than the propagation rate at early time $t$.
Appendix C Proof of Theorem 3.3
This proof follows the one proposed in Di Giovanni et al. (2023) (Appendix B).
Proof.
We proceed by induction and we show only the inductive step (i.e., $\ell>1$), since the case $\ell=1$ is straightforward (we refer to Di Giovanni et al. (2023) for more details). Assuming the Einstein summation convention and given that $\mathbf{\hat{W}}=\mathbf{W}-\mathbf{W}^{\top}$, $\mathbf{\hat{Z}}=\mathbf{Z}+\mathbf{Z}^{\top}$, and $\alpha,\beta\in[p]$, we have:
$\left|\frac{\partial\mathbf{x}_{v}^{\alpha}(\ell+1)}{\partial\mathbf{x}_{u}^{\gamma}(0)}\right|$ $=\left|\sigma^{\prime}\Bigl{(}c_{r}\mathbf{\hat{W}}^{(\ell)}_{\alpha,k}\frac{\partial\mathbf{x}_{v}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}+c_{a}\mathbf{V}^{(\ell)}_{\alpha,k}\mathbf{A}_{vz}\frac{\partial\mathbf{x}_{z}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}+\beta c_{b}\mathbf{\hat{Z}}^{(\ell)}_{\alpha,k}\mathbf{S}_{vz}\frac{\partial\mathbf{x}_{z}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}\Bigr{)}\right|$
$\leq|\sigma^{\prime}|\Bigl{(}c_{r}|\mathbf{\hat{W}}^{(\ell)}_{\alpha,k}|\left|\frac{\partial\mathbf{x}_{v}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}\right|+c_{a}|\mathbf{V}^{(\ell)}_{\alpha,k}|\mathbf{A}_{vz}\left|\frac{\partial\mathbf{x}_{z}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}\right|+\beta c_{b}|\mathbf{\hat{Z}}^{(\ell)}_{\alpha,k}|\mathbf{S}_{vz}\left|\frac{\partial\mathbf{x}_{z}^{k}(\ell)}{\partial\mathbf{x}_{u}^{\gamma}(0)}\right|\Bigr{)}$
$\leq c_{\sigma}w\Bigl{(}c_{r}\left\|\frac{\partial\mathbf{x}_{v}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|+c_{a}\mathbf{A}_{vz}\left\|\frac{\partial\mathbf{x}_{z}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|+\beta c_{b}\mathbf{S}_{vz}\left\|\frac{\partial\mathbf{x}_{z}(\ell)}{\partial\mathbf{x}_{u}(0)}\right\|\Bigr{)}$
$\leq c_{\sigma}w(c_{\sigma}wp)^{\ell}\Bigl{(}c_{r}((c_{r}\mathbf{I}+c_{a}\mathbf{A}+\beta c_{b}\mathbf{S})^{\ell})_{vu}+c_{a}\mathbf{A}_{vz}((c_{r}\mathbf{I}+c_{a}\mathbf{A}+\beta c_{b}\mathbf{S})^{\ell})_{vz}+$
$\hskip 162.18062pt+\beta c_{b}\mathbf{S}_{vz}((c_{r}\mathbf{I}+c_{a}\mathbf{A}+\beta c_{b}\mathbf{S})^{\ell})_{vz}\Bigr{)}$
$\leq c_{\sigma}w(c_{\sigma}wp)^{\ell}\Bigl{(}(c_{r}\mathbf{I}+c_{a}\mathbf{A}+\beta c_{b}\mathbf{S})^{\ell+1}\Bigr{)}_{vu}$
where ||$\cdot$|| ⋅ | denotes an absolute value of a real number, $w$ is the maximal entry-value over all weight matrices, and $c_{\sigma}$, $c_{r}$, $c_{a}$ , and $c_{b}$ are the Lipschitz maps of the components in the computation of SWAN. We can now sum over $\alpha$ on the left, generating an extra $p$ on the right side. ∎
Appendix D Datasets Details
D.1 Graph Transfer Dataset
We built the graph transfer datasets upon (Di Giovanni et al., 2023). In each task, graphs use identical topology, but, differently from the original work, nodes are initialized with random input features sampled from a uniform distribution in the interval $[0,0.5)$. In each graph, we selected a source node and target node and initialized them with labels of value “1” and “0”, respectively. We sampled graphs from three graph distributions, i.e., line, ring, and crossed-ring. Figure 4 shows a visual exemplification of the three types of graphs when the distance between the source and target nodes is 5. Specifically, ring graphs are cycles of size $n$, in which the target and source nodes are placed at a distance of $\lfloor n/2\rfloor$ from each other. Crossed-ring graphs are also cycles of size $n$, but include crosses between intermediate nodes. Even in this case, the distance between source and target nodes remains $\lfloor n/2\rfloor$. Lastly, the line graph contains a path of length $n$ between the source and target node. In our experiments, we consider a regression task, whose aim is to swap source and target node labels while maintaining intermediate nodes unchanged. We use an input dimension of 1, and the distance between source and target nodes is equal to 3, 5, 10, and 50. We generated 1000 graphs for training, 100 for validation, and 100 for testing.
(a) Line
(b) Ring
(c) Crossed-Ring
Figure 4: Line, ring, and crossed-ring graphs where the distance between source and target nodes is equal to 5. Nodes marked with “S” are source nodes, while the nodes with a “T” are target nodes.
D.2 Graph Property Prediction
In our experiments on graph property prediction, we followed the data generation procedure outlined in (Gravina et al., 2023). Graphs were randomly selected from various graph distributions, including Erdős–R’enyi, Barabasi-Albert, tree, and grid. Each graph contained between 25 and 35 nodes, where each node is assigned with random identifiers as input features sampled from a uniform distribution in the interval $[0,1)$. The target values represented single-source shortest paths, node eccentricity, and graph diameter. The dataset comprised a total of 7040 graphs, with 5120 used for training, 640 for validation, and 1280 for testing.
D.3 Long Range Graph Benchmark
In the Long Range Graph Benchmark section, we considered the “Peptides-func” and “Peptides-struct” datasets (Dwivedi et al., 2022). The graphs correspond to 1D amino acid chains, and they are derived such that the nodes correspond to the heavy (non-hydrogen) atoms of the peptides while the edges represent the bonds between them. Peptides-func is a multi-label graph classification dataset, with a total of 10 classes based on the peptide function, e.g., Antibacterial, Antiviral, cell-cell communication, and others. Peptides-struct is a multi-label graph regression dataset based on the 3D structure of the peptides. Specifically, the task consists of the prediction of the inertia of the molecules using the mass and valence of the atoms, the maximum distance between each atom-pairs, sphericity, and the average distance of all heavy atoms from the plane of best fit. Both Peptides-func and Peptides-struct consist of 15,535 graphs with a total of 2.3 million nodes. We applied stratified splitting to generate balanced train–valid–test dataset splits, using the ratio of 70%–15%–15%, respectively.
Appendix E Experimental Details
Our method is implemented in PyTorch (Paszke et al., 2017) and PyTorch-Geometric (Fey & Lenssen, 2019). All source codes will be released upon publication. Our experimental results were obtained using NVIDIA RTX-3090 and A100 GPUs.
E.1 SWAN versions
In our experiments, we explore four variants of our SWAN method, depending on the implementation of $\Phi$ and $\Psi$, as reported in Table 4. These versions can be grouped into two main groups, which differ in the resulting non-dissipative behavior. In the first group, which contains SWAN and SWAN-learn both $\Phi$ and $\Psi$ lead to purely imaginary eigenvalues, since we use symmetric graph shift operators and antisymmetric weight matrices, thus allowing for a node- and graph-wise non-dissipative behavior. The last group, which includes SWAN-ne and SWAN-learn-ne, can deviate from being globally non-dissipative. Indeed, if the weight matrix, ${\bf V}$, is an arbitrary matrix, then the eigenvalues of the Jacobian matrix of the system, $\mathbf{J}(t)$, are contained in a neighborhood of the imaginary axis with radius $r\leq||\mathbf{V}||$ (Bauer-Fike’s theorem (Bauer & Fike, 1960)). Although this result does not guarantee that the eigenvalues of the Jacobian are imaginary, in practice it crucially limits their position, limiting the dynamics of the system on the graph to show at most moderate amplification or loss of signals through the graph.
Table 4: The grid of the evaluated SWAN versions. We consider $\hat{{\bf A}}$ to be either the original adjacency matrix ${\bf A}$ or the symmetric normalized adjacency matrix, ${\bf D}^{-1/2}{\bf A}{\bf D}^{-1/2}$, while $\tilde{{\bf A}}$ is the random walk normalized adjacency matrix, ${\bf D}^{-1}{\bf A}$. The learned versions (-learn) employ learnable $\hat{{\bf A}}$ and $\tilde{{\bf A}}$ as described in Appendix A.2, here referred as $\hat{{\bf A}}_{\bf F}$ and $\tilde{{\bf A}}_{\bf F}$.
Name $\Phi$ $\Psi$
Weight Antisymmetry Only
  SWANβ=0 $(\hat{{\bf A}}+\hat{{\bf A}}^{\top})\mathbf{X}(\mathbf{V}-\mathbf{V}^{\top})$
Bounded Non-Dissipative
  SWAN-ne $\hat{{\bf A}}{\bf X}{\bf V}$ $(\tilde{{\bf A}}-\tilde{{\bf A}}^{\top})\mathbf{X}(\mathbf{Z}+\mathbf{Z}^{\top})$
  SWAN-learn-ne $\hat{{\bf A}}_{\bf F}\mathbf{X}\mathbf{V}$ $(\mathbf{\tilde{A}}_{\bf F}-\mathbf{\tilde{A}}_{\bf F}^{\top})\mathbf{X}(\mathbf{Z}+\mathbf{Z}^{\top})$
Global and Local Non-Dissipative
  SWAN $(\hat{{\bf A}}+\hat{{\bf A}}^{\top})\mathbf{X}(\mathbf{V}-\mathbf{V}^{\top})$ $(\tilde{{\bf A}}-\tilde{{\bf A}}^{\top})\mathbf{X}(\mathbf{Z}+\mathbf{Z}^{\top})$
  SWAN-learn $(\hat{{\bf A}}_{\bf F}+\hat{{\bf A}}_{\bf F}^{\top})\mathbf{X}(\mathbf{V}-\mathbf{V}^{\top})$ $(\mathbf{\tilde{A}}_{\bf F}-\mathbf{\tilde{A}}_{\bf F}^{\top})\mathbf{X}(\mathbf{Z}+\mathbf{Z}^{\top})$
E.2 Graph Transfer Task
We design each model as a combination of three main components. The first is the encoder which maps the node input features into a latent hidden space; the second is the graph convolution (i.e., SWAN or the other baselines); and the third is a readout that maps the output of the convolution into the output space. The encoder and the readout share the same architecture among all models in the experiments.
We perform hyperparameter tuning via grid search, optimizing the Mean Squared Error (MSE) computed on the node features of the whole graph. We train the models using Adam optimizer for a maximum of 2000 epochs and early stopping with patience of 100 epochs on the validation loss. For each model configuration, we perform 4 training runs with different weight initialization and report the average of the results. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment.
E.3 Graph Property Prediction
We employ the same datasets, hyperparameter space, and experimental setting presented in Gravina et al. (2023). Therefore, we perform hyperparameter tuning via grid search, optimizing the Mean Square Error (MSE), training the models using Adam optimizer for a maximum of 1500 epochs, and early stopping with patience of 100 epochs on the validation error. For each model configuration, we perform 4 training runs with different weight initialization and report the average of the results. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
E.4 Long-Range Graph Benchmarks
We employ the same datasets and experimental setting presented in Dwivedi et al. (2022). Therefore, we perform hyperparameter tuning via grid search, optimizing the Average Precision (AP) in the Peptide-func task and the Mean Absolute Error (MAE) in the Peptide-struct task, training the models using AdamW optimizer for a maximum of 300 epochs. For each model configuration, we perform 3 training runs with different weight initialization and report the average of the results. Also, we follow the guidelines in (Dwivedi et al., 2022; Gutteridge et al., 2023) and stay within the 500K parameter budget. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
E.5 Hyperparameters
In Table 5 we report the grids of hyperparameters employed in our experiments by each method. We recall that the hyperparameters $\epsilon$, $\gamma$, and $\hat{{\bf A}}$ refer to our method and ADGN, while $\beta$ only to our method. Moreover, we note that, in each graph transfer task, we use a number of layers that is equal to the distance between the source and target nodes.
Table 5: The grid of hyper-parameters employed during model selection for the graph transfer tasks (Transfer), graph property prediction tasks (GraphProp), and Long Range Graph Benchmark (LRGB). We observe that, in each graph transfer task, we use a number of layers that is equal to the distance between the source and target nodes.
Hyperparameters Values
Transfer GraphProp LRGB
Optimizer Adam Adam AdamW
Learning rate 0.001 0.003 0.001,  0.0005
Weight decay 0 $10^{-6}$ 0,  0.0001
N. layers 3, 5, 10, 50 1, 5, 10, 20 5,  8,  16,  32
embedding dim 64 10, 20, 30 64,  128
$\hat{{\bf A}}$ ${\bf A}$, ${\bf D}^{-1/2}{\bf A}{\bf D}^{-1/2}$ ${\bf A}$, ${\bf D}^{-1/2}{\bf A}{\bf D}^{-1/2}$ ${\bf A}$, ${\bf D}^{-1/2}{\bf A}{\bf D}^{-1/2}$
$\sigma$ tanh tanh tanh
$\epsilon$ 0.5, 0.1 1, 0.1, 0.01 1, 0.01
$\gamma$ 0.1 1, 0.1 1, 0.1
$\beta$ 1, 0.1, 0.01, -1 2, 1, 0.5, 0.1, -0.5, -1 1, 0.1
Appendix F Complexity and Runtimes
Complexity Analysis.
Our SWAN architecture remains within the computational complexity of MPNNs (e.g., Morris et al. (2019); Xu et al. (2019)) and other DE-GNNs such as GRAND (Chamberlain et al., 2021b). Specifically, each SWAN layer is linear in the number of nodes $|\cal{V}|$ and edges $|\cal{E}|$, therefore it has a time complexity of $\mathcal{O}(|\cal{V}|+|\cal{E}|)$. Assuming we compute $L$ steps of the ODE defined by SWAN, the overall complexity is $\mathcal{O}\left(L\cdot(|\cal{V}|+|\cal{E}|)\right)$.
Runtimes. We measure the training and inference (i.e., the epoch-time and test-set time) runtime of SWAN on the Peptides-struct dataset and compare it with other baselines. Our runtimes show that SWAN is able to obtain high performance, while retaining a linear complexity runtimes that is parallel to other MPNNs and DE-GNNs. Our measurements are reported in Table 6, and presents the training and inference runtimes in seconds, as well as the obtained metric for that task, the mean-absolute-error (MAE), for reference. The runtimes were measured on an NVIDIA RTX-3090 GPU with 24GB of memory. In all time measurements, we use a batch size of 64, 128 feature channels, and 5 layers. For reference, we also provide the reported downstream task performance of each method.
Table 6: Measured training and inference runtimes in seconds, and the obtain MAE of SWAN and other baselines on the Peptides-func dataset.
Method Training Inference MAE ($\downarrow$)
GCN 2.90 0.32 0.3496±0.0013
GraphGPS+LapPE 23.04 2.39 0.2500±0.0005
GraphCON 3.03 0.27 0.2778±0.0018
ADGN 2.83 0.25 0.2874±0.0021
SWAN 2.88 0.24 0.2571±0.0018
SWAN-learn 2.93 0.26 0.2485±0.0009
Appendix G Additional Comparisons
In Table 2 we are interested in directly comparing the performance of SWAN and its non-dissipative properties with other MPNNs and DE-GNNs, as well as the common approach of using graph transformer to address long-range interaction modeling. We now provide additional comparisons with other methods that utilize multi-hop information, and are therefore more computationally expensive than our SWAN, while also utilizing additional features such as the Laplacian positional encoding. In our evaluation of SWAN, we chose not to use additional feature enhancements, in order to provide a clear exposition of the contribution and importance of the local and global non-dissipativity offered by SWAN. The additional comparisons are given in Table 7. In the Peptide-struct task, SWAN demonstrates superior performance compared to all the compared methods. On the Peptide-func task, SWAN outperforms MPNNs, Graph Transformers, DE-GNNs, and some multi-hop GNNs performance to other methods, while being second to multi-hop methods such as DRew. The PascalVOC-SP results show the effectiveness of SWAN with other DE-GNNs (GRAND, GraphCON, ADGN), which are the focus of our paper, while offering competitive results to multi-hop and transformer methods. However, it is essential to note that multi-hop GNNs incur higher complexity, while SWAN maintains a linear complexity. Therefore, we conclude that SWAN offers a highly effective approach for tasks that require long-range interactions, as in the LRGB benchmark.
Table 7: Performance of various classical, multi-hop and static rewiring MPNN, graph Transformer benchmarks, DE-GNNs, and our SWAN across two LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded. Beseline results are reported from (Gutteridge et al., 2023).
Model Peptides- Peptides- Pascal
func struct voc-sp
AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$
MPNNs
GCN 0.5930±0.0023 0.3496±0.0013 0.1268±0.0060
GINE 0.5498±0.0079 0.3547±0.0045 0.1265±0.0076
GCNII 0.5543±0.0078 0.3471±0.0010 0.1698±0.0080
GatedGCN 0.5864±0.0077 0.3420±0.0013 0.2873±0.0219
GatedGCN+PE 0.6069±0.0035 0.3357±0.0006 0.2860±0.0085
Multi-hop GNNs
DIGL+MPNN 0.6469±0.0019 0.3173±0.0007 0.2824±0.0039
DIGL+MPNN+LapPE 0.6830±0.0026 0.2616±0.0018 0.2921±0.0038
MixHop-GCN 0.6592±0.0036 0.2921±0.0023 0.2506±0.0133
MixHop-GCN+LapPE 0.6843±0.0049 0.2614±0.0023 0.2218±0.0174
DRew-GCN 0.6996±0.0076 0.2781±0.0028 0.1848±0.0107
DRew-GCN+LapPE 0.7150±0.0044 0.2536±0.0015 0.1851±0.0092
DRew-GIN 0.6940±0.0074 0.2799±0.0016 0.2719±0.0043
DRew-GIN+LapPE 0.7126±0.0045 0.2606±0.0014 0.2692±0.0059
DRew-GatedGCN 0.6733±0.0094 0.2699±0.0018 0.3214±0.0021
DRew-GatedGCN+LapPE 0.6977±0.0026 0.2539±0.0007 0.3314±0.0024
Transformers
Transformer+LapPE 0.6326±0.0126 0.2529±0.0016 0.2694±0.0098
SAN+LapPE 0.6384±0.0121 0.2683±0.0043 0.3230±0.0039
GraphGPS+LapPE 0.6535±0.0041 0.2500±0.0005 0.3748±0.0109
DE-GNNs
GRAND 0.5789±0.0062 0.3418±0.0015 0.1918±0.0097
GraphCON 0.6022±0.0068 0.2778±0.0018 0.2108±0.0091
ADGN 0.5975±0.0044 0.2874±0.0021 0.2349±0.0054
Ours
SWAN 0.6313±0.0046 0.2571±0.0018 0.2796±0.0048
SWAN-learn 0.6751±0.0039 0.2485±0.0009 0.3192±0.0250
Appendix H Derivation of the Graph-wise Jacobian
Recall the ODE that defines SWAN in Section 3.2:
$\frac{\partial\mathbf{X}(t)}{\partial t}=\sigma\Bigl{(}\mathbf{X}(t){(\mathbf{W}-\mathbf{W}^{\top})}+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-\mathbf{V}^{\top})}+\beta{(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})}\Bigr{)}.$ (25)
The Jacobian of Equation 25 with respect to $\mathbf{X}(t)$ is:
$\mathbf{M}_{1}$ $=\sigma^{\prime}\left(\mathbf{X}(t){(\mathbf{W}-\mathbf{W}^{\top})}+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-\mathbf{V}^{\top})}+\beta{(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})}\right)$ (26)
$\mathbf{M}_{2}$ $={(\mathbf{W}-\mathbf{W}^{\top})}+{(\mathbf{V}-\mathbf{V}^{\top})^{\top}\otimes(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})}+\beta(\mathbf{Z}+\mathbf{Z}^{\top})^{\top}\otimes(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})$ (27)
where $\otimes$ is the Kronecker product. To analyze $\textbf{M}_{1}$, we use the following identity for arbitrary matrices $\mathbf{A},\mathbf{X},\mathbf{B}$ with appropriate dimensions (i.e., not related to the notations of this paper):
$\rm{vec}(\mathbf{AXB})=\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right){\rm{vec}}(\mathbf{X})$ (28)
Using the identity from Equation 28, we can rewrite $\mathbf{M}_{1}$ as follows:
$\mathbf{M}_{1}$ $=\sigma^{\prime}\Bigl{(}\mathbf{I}\mathbf{X}(\mathbf{W}-\mathbf{W}^{\top})+(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-\mathbf{V}^{\top})+\beta(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})\Bigr{)}$
$={\rm{diag}}\left(\rm{vec}(\sigma^{\prime}\Bigl{(}\mathbf{I}\mathbf{X}(\mathbf{W}-\mathbf{W}^{\top})+(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-\mathbf{V}^{\top})+\beta(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})))\right)$
$={\rm{diag}}\Bigl{(}\sigma^{\prime}\Bigl{(}((\mathbf{W}-\mathbf{W}^{\top})^{\top}\otimes\mathbf{I)}\rm{vec}(\mathbf{X})+$
$\hskip 85.35826pt+((\mathbf{V}-\mathbf{V}^{\top})^{\top}\otimes(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}))\rm{vec}(\mathbf{X}(t))+$
$\hskip 85.35826pt+\beta((\mathbf{Z}+\mathbf{Z}^{\top})^{\top}\otimes(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top}))\rm{vec}(\mathbf{X}(t))\Bigr{)}\Bigr{)},$ (29)
where $\mathbf{I}$ is the identity matrix. Therefore $\mathbf{M}_{1}$ is a diagonal matrix.
We note that Equation 25 is the result of the composite function $\sigma(g(\mathbf{X}(t)))$, where $\sigma$ is the activation function and $g(\mathbf{X}(t))=\mathbf{X}(t){(\mathbf{W}-\mathbf{W}^{\top})}+{(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})\mathbf{X}(t)(\mathbf{V}-\mathbf{V}^{\top})}+\beta{(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})\mathbf{X}(t)(\mathbf{Z}+\mathbf{Z}^{\top})}$.
Therefore, $\textbf{M}_{2}$ results from the derivative of $g$ with respect to $\mathbf{X}(t)$. Considering the $\rm{vec}$ operator, we have
$\textbf{M}_{2}$ $=\rm{vec}(g^{\prime}(\mathbf{X}(t)))$
$=g^{\prime}(\rm{vec}(\mathbf{X}(t)))$
$={(\mathbf{W}-\mathbf{W}^{\top})}+{(\mathbf{V}-\mathbf{V}^{\top})^{\top}\otimes(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top})}+\beta(\mathbf{Z}+\mathbf{Z}^{\top})^{\top}\otimes(\tilde{\mathbf{{A}}}-\tilde{\mathbf{{A}}}^{\top})$ (30)
Appendix I The Stability of the Jacobian
As discussed in Section 2.3, assuming that the Jacobian of the underlying system does not change significantly over time allows us to analyze the system from an autonomous system perspective (Ascher & Petzold, 1998) and mirrors prior approaches (Ruthotto & Haber, 2020; Chen et al., 2018a; Ruthotto & Haber, 2019; Chang et al., 2019; Chamberlain et al., 2021b; Gravina et al., 2023). In addition to building on existing literature, below, we provide an empirical measurement on a real-world dataset (pepties-func) of the Jacobian of our SWAN over time (layers). For reference, we compare it with the Jacobian of GCN. As can be seen from Figure 7, the Jacobian of SWAN has a minimal Jacobian change over time with an average of 0.6% between layers, while the change in the Jacobian over time in GCN is 40% on average.
$\frac{|\mathbf{J}^{(\ell)}-\mathbf{J}^{(\ell-1)}|}{|\mathbf{J}^{(\ell-1)|}}$
Figure 5: Arithmetic scale
$\frac{|\mathbf{J}^{(\ell)}-\mathbf{J}^{(\ell-1)}|}{|\mathbf{J}^{(\ell-1)|}}$
Figure 6: Logarithmic scale
Figure 7: The relative change in Jacobian of GCN and our SWAN over layers, measured on the peptides-func data.